# Cloud2BIM: An open-source automatic pipeline for efficient conversion of large-scale point clouds into IFC format

### v1

@@ 4.3 Speed performance analysis
This subsection presents the computational performance results for the test datasets described in Section 2. The evaluations were conducted on a Dell Precision 3591 notebook with an Intel Core Ultra 165H CPU (1.40 GHz), 32 GB RAM, a PM981 NVMe Samsung SSD, and a Windows 11 operating system. The software, developed in Python 3.10, was tested with datasets chosen for their distinctive properties, as detailed in Section 2. Table 8 summarizes the computation times alongside parameters of the resulting models and their point clouds.
@@
While Python’s interpreted nature might not traditionally align with high-performance applications compared to compiled languages like C# or C++, it offered significant advantages in terms of development speed. We leveraged data vectorization capabilities provided by NumPy arrays to boost performance, which proved particularly effective for smaller datasets. For example, the entire reconstruction of the Kladno station, which is relatively modest in scale, was completed in just 32 seconds on standard hardware specifications. In contrast, a more extensive dataset like the Prague hotel, comprising over 40 million points, required 30 minutes to process. This increase in processing time is mainly due to the complexity involved in pairing and joining numerous wall segments. This task could benefit from a compiled language implementation once the software’s development reaches maturity.
@@
Despite its current developmental state, Cloud2BIM outperforms similar software solutions, as evidenced by a comparative analysis with the work by Macher et al. (2017), who did not incorporate opening identification and have not released their code as open-source. Our software demonstrated approximately 7$\times$ the processing efficiency: in a benchmark test, our system processed a 7 million point dataset at a rate of 6.9 million points per minute, compared to the 1.25 million points per minute achieved by the aforementioned study on a 10 million-point dataset. For context, their system required 8 minutes to process a house with approximately 10 million points covering 200 m2 and 19 minutes for the main block of an office building encompassing 1000 m2 with approximately 16 million points (Table 7). Shi et al. Shi et al. (2019) also used the ETH Zürich synth3 datasetand its processing took them 359 seconds without providing the IFC output while we reached the total processing time, including IFC generation, of 60 seconds (Table 8).
@@ 4.4 Final remarks and future developments
Cloud2BIM is still evolving, with ongoing enhancements such as the integration of furniture recognition using deep-learning models (Qian et al., 2022). Furthermore, we are developing a graphical user interface that will allow users to smoothly adjust and calibrate input parameters. This interface will feature live visualization, enabling immediate observation of how changes to calibrated parameters affect the model, thus ensuring precise and user-friendly software interaction.
Given the popularity of the RANSAC algorithm for segmenting planar elements from point clouds, one might question why it was not employed in our development. The decision was based on significant concerns regarding performance and the calibration of inputs. While RANSAC is adept at managing noise and outliers, its iterative and random sampling method can inadvertently produce spurious planes that do not correspond to actual planar surfaces (Xu et al., 2016), especially in complex indoor environments. This issue is compounded by the inherent uncertainty in the algorithm’s hypothesis generation phase (Li et al., 2017). Although robust methods are designed to address these inaccuracies, they typically involve point-based operations that are inefficient for indoor settings.
Building on the capabilities of our software, we are exploring technologies to further enrich the BIM models with detailed features and materials not discernible through geometry alone. Inspired by Adán et al. (2018), who utilized RGB data from point clouds to detect minor architectural elements like electrical sockets, we plan to adapt similar image analysis techniques. Although their approach currently lacks IFC export functionality, restricting its integration with BIM tools, it provides a promising avenue for adding realistic details such as color textures and material finishes to our models. This enhancement will not only increase the richness of the visual and functional aspects of the models but also supports comprehensive documentation and precise refurbishment planning.
@@ 5 Conclusion
The segmentation and classification process within Cloud2BIM employs specialized algorithms tailored to each building element type, ensuring high accuracy and superior speed compared to other solutions. Basically, Cloud2BIM is able to segment and classify:
@@ Acknowledgement
This work was funded by the European Union under the project ROBOPROX (no. CZ.02.01.01/00/22_008/0004590), by the European Union’s Horizon Europe Framework Programme (call HORIZON-CL4-2021-TWIN-TRANSITION-01-11) under grant agreement No. 101058580, project RECONMATIC (Automated solutions for sustainable and circular construction and demolition waste management), and the Czech Technical University in Prague, grant agreement No. SGS24/003/OHK1/1T/11 (Application of state-of-the-art technologies for enhancing sustainability of the construction sector).

### v2

@@ 3.3 Speed performance analysis
This subsection presents the computational performance results for the test datasets described in Section 2.1. The evaluations were conducted on a Dell Precision 3591 notebook with an Intel Core Ultra 165H CPU (1.40 GHz), 32 GB RAM, a PM981 NVMe Samsung SSD, and a Windows 11 operating system. The software, developed in Python 3.10, was tested with datasets chosen for their distinctive properties, as detailed in Section 2.1. Table 9 summarizes the computation times alongside parameters of the resulting models and their point clouds.
@@
While Python’s interpreted nature might not traditionally align with high-performance applications compared to compiled languages like C# or C++, it offered significant advantages in terms of development speed. We leveraged data vectorization capabilities provided by NumPy arrays to boost performance, which proved particularly effective for smaller datasets. For example, the entire reconstruction of the Kladno station, which is relatively modest in scale, was completed in just 30 seconds on standard hardware specifications. In contrast, a more extensive dataset like the Prague hotel, comprising over 40 million points, required 30 minutes to process. This increase in processing time is mainly due to the complexity involved in pairing and joining numerous wall segments. This task could benefit from a compiled language implementation once the software’s development reaches maturity.
@@
Despite its current developmental state, Cloud2BIM outperforms similar software solutions, as evidenced by a comparative analysis with the work by Macher et al. [27], who did not incorporate opening identification and have not released their code as open-source. Our software demonstrated approximately 7$\times$ the processing efficiency: in a benchmark test, our system processed a 7 million point dataset at a rate of 6.9 million points per minute, compared to the 1.25 million points per minute achieved by the aforementioned study on a 10 million-point dataset. For context, their system required 8 minutes to process a house with approximately 10 million points covering 200 m2 and 19 minutes for the main block of an office building encompassing 1000 m2 with approximately 16 million points (Table 7). Shi et al. [37] also used the ETH Zürich synth3 datasetand its processing took them 359 seconds without providing the IFC output while we reached the total processing time, including IFC generation, of 60 seconds (Table 9).
3.4 Limitations
Despite the promising results achieved in this study, several limitations must be acknowledged. These limitations highlight areas for future research and potential enhancements to further improve the robustness and applicability of the proposed method.
First, the method does not currently support the detection and reconstruction of curved walls. The implemented algorithm is optimized for planar geometries, meaning that curved surfaces are either approximated by segmented planar sections or may not be processed correctly. This limitation is particularly relevant for buildings with complex architectural features, such as circular walls, domed structures, or organic geometries.
Second, the system is currently limited to detecting and reconstructing rectangular and square window and door openings. While these shapes encompass the majority of standard architectural designs, they do not account for arched, circular, or irregularly shaped openings.
Third, the method assumes a uniform height level for each floor, which may lead to inaccuracies in buildings with more complex layouts. Features such as stepped floors, mezzanines, and varying ceiling heights within the same story are not currently accounted for in the segmentation process. As a result, the accuracy of the generated BIM models may be reduced in buildings with split-level designs or other non-standard height configurations.
Finally, the accuracy and reliability of the method are highly dependent on the quality of the input point cloud. If the scanned data is noisy, sparse, or contains excessive occlusions due to furniture or other obstructions (Figure 23), the reconstruction process becomes significantly more challenging. In such cases, manual adjustments such as removal of incorrectly identified openings must be done in a BIM modeling software. To mitigate these challenges, high-quality point cloud acquisition is essential, including proper scanning techniques and preprocessing steps to filter out unwanted elements.
Figure 23: Impact of obstruction on point cloud interpretation, causing false detection of a door opening in the reconstructed IFC model.
@@ 3.5 Final remarks and future developments
Cloud2BIM is still evolving, with ongoing enhancements such as the integration of furniture recognition using deep-learning models [53]. Furthermore, we are developing a graphical user interface that will allow users to smoothly adjust and calibrate input parameters. This interface will feature live visualization, enabling immediate observation of how changes to calibrated parameters affect the model, thus ensuring precise and user-friendly software interaction.
Given the popularity of the RANSAC algorithm for segmenting planar elements from point clouds, one might question why it was not employed in our development. The decision was based on significant concerns regarding performance and the calibration of inputs. While RANSAC is adept at managing noise and outliers, its iterative and random sampling method can inadvertently produce spurious planes that do not correspond to actual planar surfaces [54], especially in complex indoor environments. This issue is compounded by the inherent uncertainty in the algorithm’s hypothesis generation phase [55]. Although robust methods are designed to address these inaccuracies, they typically involve point-based operations that are inefficient for indoor settings.
Building on the capabilities of our software, we are exploring technologies to further enrich the BIM models with detailed features and materials not discernible through geometry alone. Inspired by Adán et al. [44], who utilized RGB data from point clouds to detect minor architectural elements like electrical sockets, we plan to adapt similar image analysis techniques. Although their approach currently lacks IFC export functionality, restricting its integration with BIM tools, it provides a promising avenue for adding realistic details such as color textures and material finishes to our models. This enhancement will not only increase the richness of the visual and functional aspects of the models but also supports comprehensive documentation and precise refurbishment planning.
Moreover, considering recent advancements in autonomous mobile scanning for 3D reconstruction, as proposed by Hu et al. [56], there is potential to extend Cloud2BIM or its components to support robot-assisted autonomous scanning workflows. This would enable instant point cloud segmentation and automated 3D model reconstruction directly at the point of data capture, significantly speeding up the BIM generation process and reducing the need for post-processing.
@@ 4 Conclusion
The segmentation and classification process within Cloud2BIM employs specialized algorithms tailored to each building element type, ensuring high accuracy and efficiency; Cloud2BIM is able to segment and classify:
@@ Acknowledgement
This work was funded by the European Union under the project ROBOPROX (no. CZ.02.01.01/00/22_008/0004590), by the European Union’s Horizon Europe Framework Programme (call HORIZON-CL4-2021-TWIN-TRANSITION-01-11) under grant agreement No. 101058580, project RECONMATIC (Automated solutions for sustainable and circular construction and demolition waste management), and the Czech Technical University in Prague, grant agreement No. SGS25/005/OHK1/1T/11.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used AI language model ChatGPT 4 from OpenAI L.L.C. (CA 94110 San Francisco, USA) in order to improve readability and language of the manuscript. After using these tools, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.